OpenStudy (anonymous):

x2+8x^2+7x over x^2+15x+56. simplify

OpenStudy (lalaly):

can u rewrite the equation pweez?

OpenStudy (anonymous):

not sure how its simple do u know the answer

OpenStudy (lalaly):

is x2 >> x^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually its x^3

OpenStudy (lalaly):

x(x^2+8x+7) /(x+8)(x+7) = x(x+7)(x+1) / (x+8)(x+7) = x(x+1)/(x+8)

OpenStudy (anonymous):

Oh, I knew it was x^3. Okay, on the other question that's just like this one I posted the answer. I'll post it here too

OpenStudy (anonymous):

This is what you do. 1) x^3+8x^2+7x/x^2+15x+56 2) x(^2+8x+7)/(x+7) (x+8) 3) x(x+1)(x+7)/(x+7)(x+8) 4) x=0 : x=-7 double root : x=-8 You know that there are going to have to be three factors because of the fact that the biggest exponent value in your equation was x^3 the three indicates that there will be 3 factors.

OpenStudy (anonymous):

I hope this will suffice